find the value of k so that the line has slope m:
(k+1)x+2y=6 and m= -1/3

Answer by nerdybill(7384)

find the value of k so that the line has slope m:
(k+1)x+2y=6 and m= -1/3
.
First, take:
(k+1)x+2y=6
and manipulate it into the "slope-intercept" form:
y = mx + b
(k+1)x+2y=6
2y = -(k+1)x + 6
dividing both sides by 2:
y = (-(k+1)/2)x + 3
Now, we know the slope (in terms of k) is:
m = -(k+1)/2
.
Set the above to the given slope:
-(k+1)/2 = -1/3
(k+1)/2 = 1/3
3(k+1) = 2
(k+1) = 2/3
k = 2/3 - 1
k = 2/3 - 3/3
k = -1/3